OpenStudy (anonymous):

Solve the system of equations 10x-7y=-6 5x+7y=81

jimthompson5910 (jim_thompson5910):

What do you get when you add the equations?

OpenStudy (anonymous):

idk

jimthompson5910 (jim_thompson5910):

we get 10x-7y=-6 5x+7y=81 ---------- 15x+0y = 75 which becomes 15x = 75

jimthompson5910 (jim_thompson5910):

Do you see where to go from here?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

alright great

OpenStudy (anonymous):

dont u have 2 find y first though

jimthompson5910 (jim_thompson5910):

No, when you add the equations, the y terms will cancel out and go away.

jimthompson5910 (jim_thompson5910):

So following the method above, you solve for x first, then you use it to find y.

OpenStudy (anonymous):

okaii shanxs

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

does it end up being (5,6.28) when you put it as an ordered pair

jimthompson5910 (jim_thompson5910):

It should be (5,8)

jimthompson5910 (jim_thompson5910):

5x+7y=81 5*5+7y=81 25+7y = 81 7y = 81 - 25 7y = 56 y = 56/7 y = 8 So this shows us that the solution is (5,8)

OpenStudy (anonymous):

k shanxs again

jimthompson5910 (jim_thompson5910):

sure thing